OpenStudy (anonymous):

A block is projected at an incline at angle θ. It returns to its initial position with half of its initial speed. Show that the coefficient of kinetic friction is Uk= 3/5 tanθ

OpenStudy (anonymous):

\[PE = mg*d \sin \theta \] \[W _{f} =\mu _{k} *mg* \cos \theta * d\] \[KE _{1} = PE +W _{f}\] \[KE _{2} = PE - W _{f}\] \[\frac{ KE _{1}}{ KE _{2}} = 4 = \frac{PE + W _{f} }{PE - W _{f} }\] \[ 4 (PE - W _{f} )= PE + W _{f} \] \[\frac{ 3 }{ 5 } PE = W _{f} \] \[\frac{ 3 }{ 5 } mg∗dsinθ =μk∗mg∗\cosθ∗d\]